# Fine-Grained Perturbation Guidance via Attention Head Selection

### v3

@@ 4.1 HeadHunter: An Iterative Framework for Retrieving Effective Attention Heads
[…]
@@
Figure 5: Qualitative results of HeadHunter for style-oriented quality improvement. The first row presents the unguided result. In each subsequent row, we apply head-level guidance using additional attention heads that are incrementally selected by the HeadHunter framework. As more heads are accumulated over rounds, the generated images exhibit progressively stronger alignment with the target style while improving visual quality. Original figures and additional results conducted in FLUX.1-Dev blackforestlabs_flux1_dev can be found in Appendix D.2.
@@
6:  $\mathcal{C}\leftarrow[\,]$triangleright Initialize temporary score list
@@
9:   $\mathcal{S}_{\text{target}}\leftarrow\mathcal{S}_{\text{final}}\cup\{(l,h)\}$triangleright Define perturbation set
@@
12:   $s_{(l,h)}\leftarrow\frac{1}{M}\sum_{j=1}^{M}\mathcal{O}(\hat{x}_{j},p_{j})$triangleright Compute average objective score
@@
[…]
@@ 4.1.1 Improving general image quality
Human preference alignment perspective.
To improve image quality, various metrics can be used. A particularly interesting and increasingly popular approach leverages reward models trained on human preferences schuhmann2022laion ; kirstain2023pick ; xu2023imagereward ; zhang2024learning . Many works utilize differentiable reward models such as PickScore kirstain2023pick , HPS wu2023human , and ImageReward xu2023imagereward as supervision signals for either directly fine-tuning the model kevin2023 ; mihir2023 or adjusting the noise input to maximize the reward of the generated images eyring2024reno . However, methods that rely on the reward model’s gradients are vulnerable to reward hacking, where the model is optimized to deceive the reward model rather than genuinely improving image quality. Further, fine-tuning approaches kevin2023ddpo ; fan2023 are prone to catastrophic forgetting. On the other hand, gradient-free methods generally rely on sample-level guidance during the denoising process yeh2024demon , which requires generating and evaluating many candidate samples at each timestep. In contrast, our method identifies a set of guidance hyperparameters (attention heads) that can be reused across prompts after a single search. Although we do not directly optimize for reward scores, our approach offers a lightweight and generalizable mechanism for aligning generation with human preferences without relying on gradients or fine-tuning.
@@
We demonstrate our method using PickScore kirstain2023pick as the objective $\mathcal{O}$. HeadHunter is applied with 20 prompt-seed pairs $\mathcal{Q}$, one round ($R=1$), and a selection of $k=24$ heads per round. Additional implementation details are provided in Appendix D. Note that a single round was sufficient in our experiments to achieve substantial improvement in image quality, as the most important factor for human preference is the realism of object structures in the image, which can often be captured in a single step. In cases where users want to discover latent heads that can further improve quality but are only retrieved through multi-step searches, they may opt to use more compute by increasing the number of search rounds $R$. Qualitative results in Fig. 4 (a) show that image quality progressively improves as more HeadHunter-selected heads are included in the guidance process, supporting the compositional nature of head-level perturbations observed earlier. Fig. 4 (b) presents quantitative results in terms of FID, evaluated on 1K prompts from the MS COCO lin2014microsoft validation set. The performance improves as top-ranked heads are incrementally added, except in the high guidance scale case ($w=6.0$), where oversaturation may occur. Remarkably, using only 25% of the heads ($k=6$ out of 24) achieves comparable to or even better performance than full layer-level perturbation (shown as dotted lines for each guidance scale). These results suggest that HeadHunter can identify a compact yet effective subset of attention heads, outperforming heuristic layer-based selection strategies. Thanks to the compositionality of head-level perturbations, even a single round ($R=1$) is sufficient, as the contributions of individual heads can be combined to further improve image quality.
@@ 4.1.2 Improving style-oriented quality
While Section 4.1.1 demonstrated HeadHunter’s effectiveness in improving overall fidelity of samples, many real-world applications require more targeted control over specific visual styles, such as evoking a particular mood, or mimicking classical art techniques.
Inspired by the findings of Sec. 3.1 and Fig. 3, which reveal that certain attention heads are responsible for distinct stylistic or geometric attributes, we investigate whether HeadHunter can selectively enhance a target style through head-level guidance—while preserving or even enhancing overall image quality. We refer to this approach as style-oriented quality improvement.
@@
Fig. 6 reports quantitative metrics: PickScore kirstain2023pick and the LAION Aesthetic Score (AES) schuhmann2022laion both improve steadily as $k$ increases in the prompt-seed pairs $\mathcal{Q}$. These results confirm that HeadHunter effectively amplifies target styles in alignment with human preferences. For additional qualitative examples with SD3 (esser2024scaling, ) and FLUX.1-Dev (blackforestlabs_flux1_dev, ), we refer readers to Figs.˜21, 22, 23, 24, 25 and 26 in Appendix.
@@
Assessing generalizability to unseen content prompts.
@@ 4.1.3 Discussion
Interestingly, some heads that individually produce poor outputs can still play an essential role in composition. In Fig. 8 (a), we observe a sudden enhancement in stylistic expression (e.g., intensified red or pink hues) with inclusion of certain head set. By individually performing head-level guidance with each head in the set, we found that some head alone yields low-quality outputs but it reliably enforces a style trait. Such heads are unlikely to be selected through one-shot evaluation due to low quality, but HeadHunter’s iterative nature allows them to be integrated later as their utility emerges in composition. This highlights HeadHunter’s capacity to exploit compositional synergies among heads.
Figure 9: Layer distribution of heads selected by HeadHunter for different objectives. Effective heads are not confined to specific layers and the distribution is different with each objective, highlighting the limitations of layer-level guidance and the necessity of objective-specific head selection. (a) In general image quality setting, selections concentrate in early–mid layers (2.9% in layers 16$\geq$ 16≥ 16). (b) In style-oriented quality setting, selections spread broadly, with 24.8% in layers 16$\geq$ 16≥ 16. (c) Round-wise distributions of (b) begin with an early-layer bias similar to (a), but gradually flatten over rounds, mirroring (b). This trend aligns with qualitative observations. Round 1 primarily improves general quality, whereas stylization becomes more prominent in subsequent rounds.
Distribution of selected heads across layers.
Does the style enhancement occur even without style prompts?
Figure 11: Generated Images with linear interpolation between attention map $\mathbf{A}$ and an identity matrix $\mathbf{I}.$ While increasing $u$ enhances quality up to a point, it eventually results in over-saturation and structural over-simplification. Detailed hyperparameters can be found in Appendix F.
@@ 4.2 Controlling Guidance Intensity via Attention Map Interpolation
Although our head-level guidance provides fine-grained, objective-specific control, we propose an orthogonal approach that enables continuous modulation of the guidance effect. This allows for more precise adjustment of the perturbation intensity applied to selected heads or layers. Specifically, we introduce a simple interpolation strategy that interpolates the original attention map $\mathbf{A}$ with its perturbed counterpart. For example, the interpolated attention map for perturbed-attention guidance (PAG) ahn2024self is defined as:
@@
We present qualitative results across varying $u$ values in Fig. 11. These results show that increasing the interpolation parameter $u$ initially enhances sample quality by rectifying structural artifacts in the samples. However, beyond a certain point, it leads to oversimplified images with smoothed backgrounds and exaggerated structures. The interpolation method provides a “sweet spot” where the samples exhibit improved structure while retaining visually realistic features. By treating the identity-matrix perturbation in PAG ahn2024self as a specific point in probability distribution space, our formulation enables a smooth and principled interpolation between $\mathbf{A}$ and $\mathbf{I}$. The scalar $u\in[0,1]$ provides an intuitive and interpretable control over perturbation strength. Additionally, since this method is a simple linear interpolation of self-attention maps [^3] Although more complex paths between probability distributions (e.g., Fisher-Rao or optimal transport geodesics) exist, we find that linear interpolation—also known as a mixture-coordinate geodesic—is highly effective in practice., it can be implemented with a few lines of code. We refer to this controllable variant of PAG as Soft Perturbed-Attention Guidance (SoftPAG).
@@
Fig. 13 visualizes metric values over a grid of $(w,u)$ pairs, where brighter regions indicate better performance and red boxes denote optimal configurations. Across most metrics, the best-performing setting occurs at $u<1.0$, indicating that full replacement (i.e., $u=1$ as in PAG) is often suboptimal. These results underscore the importance of explicitly controlling perturbation strength via interpolation, rather than relying solely on the guidance scale.
[…]
@@
As running HeadHunter multiple times with different interpolation parameters $u$ is computationally expensive, it is generally sufficient to search for heads using the original PAG. Nonetheless, we can explore the effect of $u$ through a single-round search to study general quality trends. To this end, we run HeadHunter with the same setup as in Sec. 4.1.1, replacing PAG with SoftPAG (Eq. 11) and varying $u$. The results, shown in Fig. 14, report FID martin2017 scores as a function of $k$ for different $u$ values.
@@
As discussed in Sec. 3.1, the number of heads used in head-level guidance affects how strong the perturbation is. When more heads are used, the intended effect such as style enhancement can be amplified. However, this may also lead to unwanted side effects including oversaturation or oversimplification due to bias of reward models kirstain2023pick ; schuhmann2022laion . In the HeadHunter framework, heads are added step by step. This helps reinforce the target objective, but may also increase the risk of artifacts. To better control the strength of guidance, we can use SoftPAG, which allows us to mitigate the artifacts through the interpolation parameter $u$.
@@
5 Conclusion
Appendix
This appendix provides supplementary material to support the main paper. It includes:
  • Related Works
  • Additional perturbation methods and framework extensions
    • Head-level guidance using different perturbation types (Sec. 16)
    • Unified probabilistic view of attention perturbation, including SEG and temperature scaling (Sec. C)
    • Comparative metrics across different perturbation strategies (Table 2)
  • HeadHunter: Implementation and additional analysis
    • Implementation details (Sec. D)
    • Additional qualitative results and style transfer with HeadHunter (Sec. D.2)
  • Extended head analysis and compositionality
    • All-head perturbation results for SD3 and FLUX.1-Dev (Sec. E.1)
    • Interpretable heads and concept amplification (Sec. E.2E.3)
  • Experimental details and resources
    • Prompt and seed lists, model configurations, and runtime specs (Sec. F)
    • Hardware setup and computational cost discussion (Sec. F)
  • Ablation and quantitative evaluations
    • Grid search results for interpolation and guidance scale (Sec. I)
Appendix A Related Works
Diffusion Models
Attention Perturbation Guidance
Appendix B Head-level perturbation guidance with different perturbation methods
@@ C.1 Uniform Guidance
Figure 17: Generated Images with linear interpolation between attention map $\mathbf{A}$ and a uniform matrix $\mathbf{U}.$ While increasing $u$ enhances quality up to a point, it eventually results in over-saturation and structural over-simplification. Detailed hyperparameters can be found in Appendix F.
From an information-theoretic perspective, the identity matrix attention map is a special case where each row of the attention map corresponds to a Dirac delta distribution, which is the lowest possible entropy case. At the opposite extreme of the entropy spectrum lies the uniform distribution $\mathbf{U}$, whose rows are all $[\tfrac{1}{N},\dots,\tfrac{1}{N}]$; it attains the maximum entropy, allowing each query to attend equally to every key. Replacing $\mathbf{A}$ by $\mathbf{U}$ in selected heads gives Uniform Guidance (UG). Similar to the interpolation with the identity matrix in SoftPAG, we can define a distributional perturbation between the original attention matrix $\mathbf{A}$ and $\mathbf{U}$ via linear interpolation:
@@ D.2.1 Additional results
We present additional qualitative results for style-oriented quality improvement using HeadHunter in Figs.˜21, 23 and 22. Each row corresponds to the result obtained by perturbing the head set selected at round $r$ of Alg.1. The top row shows unguided generation, and the lower rows show results with an increasing number of selected heads applied. As more heads are progressively added, the generated images align better with the target style while maintaining visual plausibility.
@@
HeadHunter can benefit from with model size. Since larger models have more expressive attention heads, they can be more effectively leveraged in head-level perturbation guidance. In Fig. 21-26, we provide qualitative results of style-oriented quality improvement conducted on FLUX.1-dev [37].

### v4

@@ 4.1 HeadHunter: An Iterative Framework for Retrieving Effective Attention Heads
To address this, we introduce HeadHunter, a framework that optimizes an arbitrary objective function by iteratively selecting a subset of attention heads to perturb. Each round consists of three stages: generation, evaluation, and expansion. In the generation stage, the framework perturbs attention maps for each candidate head $(l,h)\in\mathcal{S}$ and generates samples using attention perturbation guidance with multiple prompt–seed pairs $\mathcal{Q}=\{(p_{1},s_{1}),\dots,(p_{M},s_{M})\}$. During the evaluation stage, it computes the average objective score using the user-given objective function $\mathcal{O}$ over the generated samples for each candidate. In the expansion stage, the top-$k$ performing heads are added to the selected head set $\mathcal{S}_{\text{final}}$. This process is repeated for a fixed number of rounds $R$. The full procedure is described in Algorithm 1. We validate HeadHunter both quantitatively and qualitatively across tasks such as general quality and style enhancement.
@@
6:  $\mathcal{C}\leftarrow[\,]$ $\triangleright$ Initialize temporary score list
@@
9:   $\mathcal{S}_{\text{target}}\leftarrow\mathcal{S}_{\text{final}}\cup\{(l,h)\}$ $\triangleright$ Define perturbation set
@@
12:   $s_{(l,h)}\leftarrow\frac{1}{M}\sum_{j=1}^{M}\mathcal{O}(\hat{x}_{j},p_{j})$ $\triangleright$ Compute average objective score
@@
17:  $\mathcal{S}_{\text{new}}\leftarrow$ top-$k$ heads from $\mathcal{C}$
@@
Figure 4: Results of HeadHunter for general image quality improvement. Performance improves as more top-ranked heads are added, demonstrating the effectiveness of compositional head selection via HeadHunter. The dotted horizontal line in (b) indicates the best score achieved by layer-level guidance, which is surpassed by a compact set of top-$k$ heads for $k<10$. Dashed lines indicate the FID of the best-performing layer-level perturbation for each guidance scale $w$ in Eq. 5.
@@ 4.1.1 Improving general image quality
Various metrics can be used to assess image quality, including vision-language models and learned reward functions. Following recent trends in diffusion models that optimize image preference scores [49, 27, 62, 65], we demonstrate our method using PickScore [27] as the objective $\mathcal{O}$. HeadHunter is applied with 20 prompt-seed pairs $\mathcal{Q}$, one round ($R=1$), and a selection of $k=24$ heads per round. Additional implementation details are provided in Appendix D. Qualitative results in Fig. 4 (a) show that image quality progressively improves as more HeadHunter-selected heads are included in the guidance process, supporting the compositional nature of head-level perturbations observed earlier. Fig. 4 (b) presents quantitative results in terms of FID, evaluated on 1K prompts from the MS COCO [33] validation set. The performance improves as top-ranked heads are incrementally added, except in the high guidance scale case ($w=6.0$), where oversaturation may occur. Remarkably, using only 25% of the heads ($k=6$ out of 24) achieves comparable to or even better performance than full layer-level perturbation (shown as dotted lines for each guidance scale) and can be boosted much further by using additional heads. These results suggest that HeadHunter can identify a compact yet effective subset of attention heads, outperforming heuristic layer-based selection strategies.
Figure 5: Qualitative results of HeadHunter for style-oriented quality improvement. The first row presents the unguided result. In each subsequent row, we apply head-level guidance using additional attention heads that are incrementally selected by the HeadHunter framework. As more heads are accumulated over rounds, the generated images exhibit progressively stronger alignment with the target style while improving visual quality. Original figures and additional results conducted in FLUX.1-Dev [29] can be found in Appendix D.2.
@@ 4.1.2 Improving style-oriented quality
While Section 4.1.1 demonstrated HeadHunter’s effectiveness in improving overall fidelity of samples, many real-world applications require more targeted control over specific visual styles, such as evoking a particular mood, or mimicking classical art techniques. Inspired by the findings of Sec. 3.1 and Fig. 3, which reveal that certain attention heads are responsible for distinct stylistic or geometric attributes, we investigate whether HeadHunter can selectively enhance a target style through head-level guidance—while preserving or even enhancing overall image quality. We refer to this approach as style-oriented quality improvement.
@@
Fig. 6 reports quantitative metrics: PickScore [27] and the LAION Aesthetic Score (AES) [49] both improve steadily as $k$ increases in the prompt-seed pairs $\mathcal{Q}$. These results confirm that HeadHunter effectively amplifies target styles in alignment with human preferences. For additional qualitative examples with SD3 [14] and FLUX.1-Dev [29], we refer readers to Figs.˜24, 25, 26, 27, 28 and 29 in Appendix.
Assessing generalizability to unseen content prompts.
@@
Figure 8: Role of individually weak heads. For sudden stylistic transition moments (e.g., adding global warmth in (a)), we visualize the effect of newly added heads in (b). One of these heads generate blurry reddish outputs when used alone (see red box). Although they fail to produce meaningful content when used independently, they contribute effectively when composed with previously selected structural heads. This highlights the importance of iterative strategy since those heads are unlikely to be selected by one-shot evaluation due to low quality. Additional results can be found in Fig. 21
@@ 4.1.3 Discussion
Interestingly, some heads that individually produce poor outputs can still play an essential role in composition. In Fig. 8 (a), we observe a sudden enhancement in stylistic expression (e.g., intensified red) with inclusion of certain head set. By individually performing head-level guidance with each head in the set, we found that some head alone yields low-quality outputs but it reliably enforces a style trait. Such heads are unlikely to be selected through one-shot evaluation due to low quality, but HeadHunter’s iterative nature allows them to be integrated later as their utility emerges in composition. This highlights HeadHunter’s capacity to exploit compositional synergies among heads.
@@ 4.2 Controlling Guidance Intensity via Attention Map Interpolation
Although our head-level guidance provides fine-grained, objective-specific control, we propose an orthogonal approach that enables continuous modulation of the guidance effect. This allows for more precise adjustment of the perturbation intensity applied to selected heads or layers. Specifically, we introduce a simple interpolation strategy that interpolates the original attention map $\mathbf{A}$ with its perturbed counterpart. For example, the interpolated attention map for perturbed-attention guidance (PAG) [1] is defined as:
@@
We present qualitative results across varying $u$ values in Fig. 9. These results show that increasing the interpolation parameter $u$ initially enhances sample quality by rectifying structural artifacts in the samples. However, beyond a certain point, it leads to oversimplified images with smoothed backgrounds and exaggerated structures. The interpolation method provides a “sweet spot” where the samples exhibit improved structure while retaining visually realistic features. By treating the identity-matrix perturbation in PAG [1] as a specific point in probability distribution space, our formulation enables a smooth and principled interpolation between $\mathbf{A}$ and $\mathbf{I}$. The scalar $u\in[0,1]$ provides an intuitive and interpretable control over perturbation strength. Additionally, since this method is a simple linear interpolation of self-attention maps , it can be implemented with a few lines of code. We refer to this controllable variant of PAG as Soft Perturbed-Attention Guidance (SoftPAG).
Note that this interpolation approach represents a general principle that can be applied to modulate the strength of various attention-based guidance methods, including existing ones such as Smoothed Energy Guidance [22] and interpolations toward a constant matrix (i.e., uniform attention weights). A broader discussion of the unified perspective on attention perturbation and corresponding results is provided in Appendix C.
Figure 9: Generated Images with linear interpolation between attention map $\mathbf{A}$ and an identity matrix $\mathbf{I}.$ While increasing $u$ enhances quality up to a point, it eventually results in over-saturation and structural over-simplification. Detailed hyperparameters can be found in Appendix G.
@@
Fig. 10 in Appendix visualizes metric values over a grid of $(w,u)$ pairs, where brighter regions indicate better performance and red boxes denote optimal configurations. Across most metrics, the best-performing setting occurs at $u<1.0$, indicating that full replacement (i.e., $u=1$ as in PAG) is often suboptimal. These results underscore the importance of explicitly controlling perturbation strength via interpolation, rather than relying solely on the guidance scale. Due to space limitations, we defer a detailed discussion of the effectiveness of SoftPAG and the selection of $u$ and $w$ to Appendix B.
5 Conclusion
Acknowledgments and Disclosure of Funding
This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, RS-2024-00509279, RS-2025-II212068, RS-2023-00227592, RS-202502214479, RS-2024-00457882, RS-2025-25441838, RS-2025-25441838, RS-2025-02214479, RS-2025-02217259) and the Culture, Sports, and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism (RS-2024-00345025, RS-2024-00333068, RS-2023-00222280, RS-2023-00266509), and National Research Foundation of Korea (RS-2024-00346597).
Appendix
This appendix provides supplementary material to support the main paper.
Contents
Appendix A Related Works
Diffusion models.
Attention perturbation guidance.
Appendix B Continuous control on perturbation strength via attention map interpolation
Figure 11: Effect of interpolation parameter $u$ on FID across PickScore top-$k$ head selections. As $u$ decreases, the perturbation becomes milder, shifting the optimal point (lowest FID) toward larger $k$ (i.e., more heads are needed). Notably, $u=0.5$ yields the best FID (indicated by the red dashed line), highlighting the benefit of moderate perturbation strength.
@@ Appendix B Continuous control on perturbation strength via attention map interpolation
As running HeadHunter multiple times with different interpolation parameters $u$ is computationally expensive, it is generally sufficient to search for heads using the original PAG. Nonetheless, we can explore the effect of $u$ through a single-round search to study general quality trends. To this end, we run HeadHunter with the same setup as in Sec. 4.1.1, replacing PAG with SoftPAG (Eq. 7) and varying $u$. The results, shown in Fig. 11, report FID [17] scores as a function of $k$ for different $u$ values.
@@
As discussed in Sec. 3.1, the number of heads used in head-level guidance affects how strong the perturbation is. When more heads are used, the intended effect such as style enhancement can be amplified. However, this may also lead to unwanted side effects including oversaturation or oversimplification due to bias of reward models [27, 49]. In the HeadHunter framework, heads are added step by step. This helps reinforce the target objective, but may also increase the risk of artifacts. To better control the strength of guidance, we can use SoftPAG, which allows us to mitigate the artifacts through the interpolation parameter $u$.
@@ C.1 Uniform Guidance
Figure 14: Generated Images with linear interpolation between attention map $\mathbf{A}$ and a uniform matrix $\mathbf{U}.$ While increasing $u$ enhances quality up to a point, it eventually results in over-saturation and structural over-simplification. Detailed hyperparameters can be found in Appendix G.
From an information-theoretic perspective, the identity matrix attention map is a special case where each row of the attention map corresponds to a Dirac delta distribution, which is the lowest possible entropy case. At the opposite extreme of the entropy spectrum lies the uniform distribution $\mathbf{U}$, whose rows are all $\tfrac{1}{N},\dots,\tfrac{1}{N}]$; it attains the maximum entropy, allowing each query to attend equally to every key. Replacing $\mathbf{A}$ by $\mathbf{U}$ in selected heads gives Uniform Guidance (UG). Similar to the interpolation with the identity matrix in SoftPAG, we can define a distributional perturbation between the original attention matrix $\mathbf{A}$ and $\mathbf{U}$ via linear interpolation:
@@ C.3 Softmax temperature scaling
More quantitative results.
Figure 19: Quantitative results of interpolation in SoftPAG, on layer $7,8,9,10$ with different metrics.
Figure 20: Quantitative results of interpolation in UG, on layer $7,8,9,10$ with different metrics.
@@ D.2.1 Additional results
We present additional qualitative results for style-oriented quality improvement using HeadHunter in Figs.˜24, 26 and 25. Each row corresponds to the result obtained by perturbing the head set selected at round $r$ of Alg.1. The top row shows unguided generation, and the lower rows show results with an increasing number of selected heads applied. As more heads are progressively added, the generated images align better with the target style while maintaining visual plausibility.
@@
HeadHunter can benefit from with model size. Since larger models have more expressive attention heads, they can be more effectively leveraged in head-level perturbation guidance. In Fig. 24-29, we provide qualitative results of style-oriented quality improvement conducted on FLUX.1-dev [29].
D.2.2 Additional analysis
Distribution of selected heads across layers.
Figure 22: Layer distribution of heads selected by HeadHunter for different objectives. Effective heads are not confined to specific layers and the distribution is different with each objective, highlighting the limitations of layer‐level guidance and the necessity of objective‐specific head selection. (a) In general image quality setting, selections concentrate in early–mid layers (2.9% in layers $\geq 16$). (b) In style‐oriented quality setting, selections spread broadly, with 24.8% in layers $\geq 16$. (c) Round-wise distributions of (b) begin with an early-layer bias similar to (a), but gradually flatten over rounds, mirroring (b). This trend aligns with qualitative observations. Round 1 primarily improves general quality, whereas stylization becomes more prominent in subsequent rounds.
Does the style enhancement occur even without style prompts?
Computational efficiency.
[…]